OpenStudy (anonymous):

Choose the correct simplification of (5a^6b^4c^3)^0(ac^2)^3.

OpenStudy (unklerhaukus):

\[ (5a^6b^4c^3)^0(ac^2)^3\] use these \[(n^x)^y=n^{xy}\qquad n^0=1\]

OpenStudy (anonymous):

i'm confused because some people say everything to the 0 power is 0 but others say it's 1??

OpenStudy (anonymous):

but is the answer 5a^9b^4c^9??

OpenStudy (unklerhaukus):

no thats not right,

OpenStudy (anonymous):

i'm not sure how to get the right answer then

OpenStudy (unklerhaukus):

\[(5a^6b^4c^3)^0=1\]

OpenStudy (anonymous):

ok so what i did was (5a^6b^4c^3)(a^3c^6) and got the answer 5a^9b^4c^9

OpenStudy (anonymous):

ohhh okay if it equals 1 then is the answer a^3c^6??

OpenStudy (unklerhaukus):

yeah thats right, \((5a^6b^4c^3)^0(ac^2)^3=a^3c^6 \checkmark\) you would get your first answer if the question was \[ (5a^6b^4c^3)^1(ac^2)^3=5a^9b^4c^9\]

OpenStudy (anonymous):

ok thankyou!